OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -9 - 3t. Find the instantaneous velocity at t = 8 by finding the derivative.

OpenStudy (danjs):

\[\frac{ d }{ dt }s(t) = v(t)\]

OpenStudy (danjs):

Or you can see that s(t) is linear, and the velocity is constant at all times, the slope of the line

OpenStudy (anonymous):

so how would i plug this in @DanJS

OpenStudy (danjs):

Show the derivative of s, \[v(t) = \frac{ d }{ dt }s(t) = \frac{ d }{ dt }[-9-3t]\]